dx( y sin2x) = [1+y^2+cos^x]dy
1
Expert's answer
2020-06-15T17:09:02-0400

"dx(y\sin2x)=(1+y^2+\cos^2 x)dy\\
y\sin2xdx-(1+y^2+\cos^2 x)dy=0\\
P=y\sin 2x\\
Q=-(1+y^2+\cos^2 x)\\
\frac{\partial F}{\partial y}=\sin2x\\
\frac{\partial Q}{\partial x}=-2\cos x(-\sin x)=\sin 2x"

Then "\exist" function "u(x,y)" :

"\frac{\partial u}{\partial x}=P=y\sin 2x\\
\frac{\partial u}{\partial y}=Q=-(1+y^2+\cos^2 x)\\
u(x,y)=\int y\sin2xdx=-\frac{1}{2}y\cos2x+\psi(y)\\
\frac{\partial u}{\partial y}=-\frac{1}{2}\cos2x+\psi'(y)=\\
=-\frac{1}{2}(2\cos^2x-1)+\psi'(y)=\\
=-\cos^2x+0.5+\psi'(y)\\
-\cos^2x+0.5+\psi'(y)=-1-y^2-\cos^2x\\
\psi'(y)=-1.5-y^2\\
\psi(y)=-1.5y-\frac{y^3}{3}"

Then

"u(x,y)=-0.5y\cos2x-1.5y-\frac{y^3}{3}"

solution of equation is

"-0.5y\cos2x-1.5y-\frac{y^3}{3}=c"